OpenStudy (crashonce):

If the time to travel a certain distance is needed to decrease by 20%, the speed of travel should increase by how much? @mathmath333

OpenStudy (crashonce):

@perl

OpenStudy (perl):

you can use the equation distance = rate * time

OpenStudy (perl):

since distance does not change

OpenStudy (crashonce):

i got 25%

OpenStudy (perl):

d = r*t you are given that you reduce the time by 20% d = r * ( 1+ k) * t *(1-.2) since d = d r*t = r * ( 1 + k ) * t * ( 1 - .2 )

OpenStudy (crashonce):

is it 25%

OpenStudy (perl):

so you can simplify 1 = (1 + k ) ( 1-.2)

OpenStudy (perl):

I get 25% as well

OpenStudy (crashonce):

thanks just confirming, got a few more

OpenStudy (perl):

did you use similar logic>

OpenStudy (crashonce):

ye i made set values